Weather Report for March 21, 2015

Continued high humidity and clouds again prevented the dome from being opened at the beginning of the night. Conditions rapidly improved, but as an added twist, a major technical problem surfaced with the SPOL CCD since the previous night. The problem was found to be in one of the cables that connects the CCD controller with the data-taking computer, but this took awhile to track down. By the time the detector was ready for use, the sky had mostly cleared and the humidity was just under the 90% limit for observing. Observations started at about 22:45 local time. Only a few thin clouds marred the sky from then until morning twilight and the humidity fell to under 80% for the first time in nearly 4 days. The seeing was good at between 1 and 1.5 arcsec.

This concludes the first half of Campaign 70. There will be a one-night break for an unrelated EPO observing program and then SPOL will be remounted on Sunday for the second five-night portion of the monitoring campaign.

The usual source for the infrared weather images is not currently operating, so we are left with images from another site until it is fixed.